Tris buffers are commonly used in biochemistry labs because they buffer within the physiological range of 7.1 to 9.1 due to a pKa of 8.1. To demonstrate the buffering capacity of Tris buffer, your biochemistry lab teaching assistant has given you one liter of a 0.1 M Tris buffer at pH 7.4. Add 2 mL of 1M HCl to this buffer and calculate what the new pH will be.
Answer #1

Use Henderson Hasselbalch equation to determine the ratio of conjugate base to weak acid of the original solution.

7.4 = 8.1 + log [A-]/[HA]

-0.7= log [A-]/[HA]

Taking antilog

0.20/1 =[A-]/[HA]

0.20/1.20 = 0.17 =% of buffer A-= 0.17 (0.1M) =0.017M A-

1.0/1.20 = 0.83 = % of buffer HA = 0.83(0.1M) = 0.083 M HA

determine the [H+] added based on the amount and conc. of the HCl.

[HCl] = [H+] = (0.002 L) ×(1.0 M)/(1.002 L) = 0.002 M

New [A-] = 0.017 M – 0.002 M = 0.015M A-

New [HA] = 0.083 M + 0.002 M = 0.085M HA

Using the Henderson-Hasselbalch equation calculate the new pH.

pH = 8.1 + log (0.015)/0.08 = 7.35

So 7.35 is the new pH